### v1

@@ 3.2 Spatial Mapping of NSDoH
The MBMM profile assignments for each census tract were recorded and linked with census tract geographic data. We developed a web-based interactive map using Shiny to display the MBMM results, available at https://mhn38j-carmen-rodriguez.shinyapps.io/nsdoh_profiles_app_urban/. In this map, we also added a layer delineating 2020 Urban Boundaries obtained from the Massachusetts Bureau of Geographic Information Data Hub [https://gis.data.mass.gov] to further describe the NSDoH profiles.
@@
Neighborhoods that were assigned to Profile 1 were found predominantly in surrounding areas outside of Boston (e.g. Middlesex and Essex counties). Examining the overlaid urban boundaries reveals that most census tracts in this profile are located within urbanized areas. Additionally, the average proportion of Non-Hispanic White individuals in these census tracts was 85.9%, which further corroborates the results of our model for social and community context indicators. Neighborhoods assigned to Profile 2 were identified in areas with large immigrant populations (e.g., Lawrence, Lynn, Chelsea, Springfield) and a high average proportion of Hispanic/Latino residents. Neighborhoods in Profile 3 were primarily in areas with a known university presence and an average median age of 35 years old (e.g., Cambridge, Amherst, Brookline). Neighborhoods in Profile 4 were primarily in areas south and west of Boston (e.g. Berkshires, Plymouth). Neighborhoods in Profile 5 (e.g. Saugus, Winchendon, Brockton) were close in proximity to neighborhoods in Profile 2, and they shared similar age and race and ethnicity distributions in their population. Neighborhoods in profile 6 were primarily in and immediately surrounding Boston (e.g., Quincy, Waltham, Somerville). Neighborhoods in Profile 7 were mostly in Western Massachusetts and the Cape Cod area (e.g. Franklin County, Provincetown). Neighborhoods in Profile 8 were disbursed in small pockets throughout the state (e.g., Peabody, Hyde Park, Tisbury, Adams). See Supplementary Table A.1 in the Supplementary Materials for more information about the census tracts population age and race-ethnicity distribution by NSDoH profile.
Figure 1: MBMM derived NSDoH profiles. Bars indicate the estimated posterior probability of high exposure to the NSDoH variable for a neighborhood, given assignment to an NSDoH profile. SDoH variables are ordered/colored by the NSDoH thematic domain. Abbreviations: high school(HS), EN(English), Non-Hispanic (NH), Supplemental Nutrition Assistance Program (SNAP).
@@
Table 1: Distribution of selected neighborhood social determinants of health and ACS 2015-2019 table identification.
NSDoH ACS Variable (ACS Table) Median estimate for MA
% of owner-occupied housing (DP04) 66.15
@@
% Crowding in household (DP04)  [^1] Occupied housing units with 1.01 to 1.50 and 1.51 or more occupants per room/ All occupied housing units for the same calendar year[46]. 1.28
% Occupied housing units without complete plumbing (DP04) [^2] We binarized this variable differently because the median was 0. Therefore, census tracts for which there were no households lacking complete plumbing were categorized as ”0”, and the remaining tracts were categorized as ”1” to indicate lack of plumbing. 0
Estimate median household income in the past 12 months (inflation-adjusted; B19013) 82265
@@
% Below 150% poverty level (C17002)  [^3] People living below 150% the poverty level/ All people for the same calendar year[46, 47]. 12.5
@@
% Population aged 25 years or older with less than 9 years of education (DP02) 2.5
% High School degree or higher (DP02) 93.5
% Bachelor’s degree or higher (DP02) 40.5
% Language other than English: Speak English less than ”very well”(DP02) 5.8
@@
3.3 NSDoH Profiles and Receipt of Optimal Care for EC
Table B.1 shows sociodemographic and clinical information of the analytical sample of EC patients, who received optimal care or NCCN guidelines adherent treatment (82.3%). These patients were predominantly non-Hispanic White (87.2%), foreign-born (49.7%), aged 50-64 (46.1%), privately insured (46.8%), diagnosed at stage I(91.5%), and with grade 1 tumors (52.9%). Patients were initially treated at large medical facilities (75.7%), academic medical centers (38.7%), and diagnosed by physicians specializing in family/internal medicine (47%).
Table 2 details the distribution of patient demographics based on which NSDoH profile they resided at EC diagnosis. Race-ethnicity, birthplace, insurance status at diagnosis, and initial point of care facility characteristics were significantly associated with NSDoH profiles. For example, among patients residing in neighborhoods belonging to NSDoH profile pattern 2, 45.8% were foreign-born, 59.41% had Medicare or public insurance, and 39.4% were initially treated at academic medical centers.
Figure 2 summarizes the results of the regression analysis. Although not statistically significant, compared to patients who resided in neighborhoods belonging to the NSDoH Profile 1, patients in the NSDoH Profile 2 have lower odds [OR = 0.77, 95% Credible Interval (0.56,1.07)] of receiving optimal care after adjusting for year of diagnosis, age at diagnosis, insurance status at diagnosis, and initial type of care facility.
We also examined the relationship between NSDoH profiles and the type of initial care facility (academic medical center versus other facilities). The results, adjusted for year, age, and insurance status at diagnosis, are detailed in Supplementary Materials B. Compared to profile 1, patients who resided in NSDoH profiles 3 and 6 were significantly associated with higher odds of receiving care at an academic medical center, while those in NSDoH profile 5 had lower odds.
Figure 2: Adjusted odds of receiving optimal care by NSDoH profiles (Ref= NSDoH Profile 1) for patients diagnosed with EC from 2015-2019. Bars show 95% credible intervals.
4 Discussion
We used data from the ACS to derive NSDoH profiles using a fully Bayesian multivariate Bernoulli mixture model. To our knowledge, this is the first study to apply Bayesian mixture models to neighborhood-level data in this context. This approach allowed us to estimate the number of profiles directly from the data, eliminating the need for post-hoc testing. Our model identified eight (8) NSDoH profiles which we characterized using neighborhood SDoH variables belonging to four thematic domains: household conditions and resources, economic security, educational attainment, and social and community context.
We used these profiles to conduct a regression analysis where we examined the association of NSDoH profiles and optimal care for EC. Our findings, although not statistically significant, suggest disparities in EC care. Specifically, patients residing in the profile characterized as the most disadvantaged, based on the probability distribution of SDoH variables, had lower odds of receiving optimal care. These findings are comparable to Rodriguez et al. where they examined the association between neighborhood socioeconomic status (NSES) using the Yost Index in the Surveillance, Epidemiology, and End Results (SEER) database from 2006-2015 and adherence to the NCCN guidelines[12]. Rodriguez et al. was the first study, to our knowledge, to find a decreasing trend in adherence to NCCN guidelines for EC by decreasing level of NSES. Similarly to our findings, they observed that patients residing in the most disadvantaged neighborhoods with the lowest NSES had 0.73 times lower odds of receiving NCCN guideline-concordant care compared to those in the highest NSES group. In their study, 59.5% of patients received treatment adhering to NCCN guidelines, with the lowest adherence seen among Black, Latina, and American Indian/Alaska Native women (57.1%, 54.5%, and 52.7%, respectively). These proportions are lower than those in our study (82.3%) but comparable to a study using the Women’s Health Initiative (WHI), where they found 80% of patients received NCCN adherent treatment[48]. Our results focused specifically on Massachusetts, while their analysis covered the entire United States. Additionally, similar to the WHI study, our study population was predominantly Non-Hispanic White, which may have limited our ability to analyze disparities in optimal care among racial and ethnic minority groups. This is important, as differences in standard treatment regimens may exist for some groups due to variations in the histologic subtypes and stages of endometrial cancer they present with. [49] Lastly, beyond the limitations of retrospective observational studies, the MCR lacks information on comorbidities, which may also impact differences in treatment adherence and recommendations and may explain non-adherence rates among race-ethnic minority groups.
Rodriguez et al., along with other studies examining the impact of neighborhood socioeconomic status (SES) on cancer outcomes, have used the Yost Index from the SEER database [12, 11, 13, 50]. Initially developed in 2001 to examine breast cancer incidence in California, this measure of neighborhood disadvantage, although it uses a subset of our variables, provides different information and may not be reflective of healthcare access in the current socioeconomic context. This is because the Yost Index emphasizes monetary indicators and is derived via a one-component PCA. This method uses weights or the component loadings that explain the most variance within the data to create an uncorrelated linear combination of the SES indicator variables [51]. Although PCA is a powerful and widely used method to reduce the dimensionality of high-dimensional datasets, like in the case of neighborhood data, the method suffers from interpretability issues because it outputs a linear combination of the original indicators, and therefore we cannot generalize about the original variables [51]. In contrast, our approach using finite mixture models provides a holistic way to identify patterns within a dataset by modeling the data as a mixture of Bernoulli distributions. This approach provides interpretable clusters of neighborhoods making it easier to understand the heterogeneity of neighborhood data, as well as the defining characteristics of each profile. We used a Bayesian framework for mixture models because it provides the flexibility of estimating model parameters and the number of profiles simultaneously. Specifically, our approach treats the number of profiles like any other unknown parameter and puts a prior distribution on it, which provides a robust method for consistently estimating the number of profiles [52, 44, 53]. Utilizing a Bayesian framework to estimate our model allowed us to borrow information within and across other neighborhoods to generate these profiles and provide a more comprehensive view of how these factors interrelate with one another at the neighborhood level. Overall, using our NSDoH profiles to assess association with health outcomes, such as receipt of optimal care for EC, may give a more nuanced view of how SDoH work overlap and co-occur within communities rather than in isolation.
This model is still met with limitations. First, model-based clustering is reflective of the data we use. In this study, we used 5-year estimates from the ACS from 2015-2019 for MA. Thus, the current NSDoH profiles may vary across different ACS survey waves due to changes in the distribution of SDoH variables, as well as across different geographies. Our binarized thresholds were based on median values for the state of Massachusetts. A different state would yield different cutoffs and, ultimately, different profiles. Second, we made the decision to dichotomize the SDoH variables based on the skewed distributions found in the 2015-2019 cycles. This can sometimes result in a slight loss of information and reduced precision of the profile estimates since the full distribution of these variables was not considered. Future work should explore the development and implementation of a fully Bayesian multivariate beta mixture model that can flexibly accommodate bounded data, such as what we see in census-level data. Third, our analysis used aggregate data from the Census to construct NSDoH profiles. These variables are reflective of population-level characteristics rather than individual-level data. To avoid ecological fallacies, relationships observed at the aggregate level cannot be assumed to hold for individual EC patients residing within these neighborhoods. Lastly, consistent with other similar studies, our analysis is cross-sectional [23] and does not account for the changes in neighborhood demographics over time as the result of fair housing policies and gentrification. We focused our time frame on 2015-2019, which was based on the ACS 5-year estimates of data. Future studies should incorporate multiple survey cycles to assess how demographic shifts influence NSDoH profiles and, consequently, the access to and quality of healthcare and other outcomes for their residents.
5 Conclusion
Bayesian mixture models offer promising applications for neighborhood-level data. Our approach enabled us to characterize and assign NSDoH profile patterns to census tracts in Massachusetts. Geospatial mapping of NSDoH profiles demonstrated how we can leverage these tools to identify areas for targeted interventions. The profiles derived for the model were applied to examine the association with receipt of optimal care for endometrial cancer patients of certain NSDoH profiles. However, these profiles are translatable to a myriad of other outcomes and exploratory analyses.
Abbreviations
SDoH: Social determinants of health
NSDoH: Neighborhood social determinants of health
EC: Endometrial cancer
ACS: American Community Survey
MA: Massachusets
ADI: Area Deprivation Index
PCA: Principal Components Analysis
SES: Socioeconomic Status
FMM:Finite Mixture Models
LCA: Latent Class Analysis
LPA: Latent Profile Analysis
MBMM: Multivariate Bernoulli Mixture Model
SNAP: Supplemental Nutrition Assistance Program
GED: General Education Development
MC3: Metropolis-coupled Markov chain Monte Carlo algorithm
MCR: Massachusetts Cancer Registry
MDPH: Massachusetts Department of Public Health
NCCN: National Comprehensive Cancer Network
Declarations
Funding
This project was supported by the Harvard Data Science Initiative Bias2 program. Dr. Stephenson is supported by NHLBI 1K01HL166442-01. Carmen Rodriguez Cabrera is supported by NIH T32CA009337. The funders had no role in the conceptualization, design, data analysis, decision to publish, or preparation of the manuscript.
Competing interests
The authors declare that they have no competing interests.
Ethics approval and consent to participate
The institutional review board at the Massachusetts Department of Public Health (MDPH) approved the ethical clearance for the MCR-MDPH data in partnership with Harvard University School of Public Health.
Consent for publication
Not applicable.
Availability of data and materials
The R code used for this manuscript has been made publicly available in GitHub, https://github.com/cbrodriguez01/ecbayesbinmix. This also includes all the American Community Survey data we used for multivariate Bernoulli mixture model, which is also publicly available at https://www.census.gov/programs-surveys/acs/data.html. The data that support the findings from the regression analysis in this study are available from the Massachusetts Cancer Registry-Massachusetts Department of Public Health (MCR-MDPH), but restrictions apply to the availability MCR data, which were used under license for the current study, and so are not publicly available. Data are, however, available from the authors upon reasonable request and with permission of MCR-MDPH.
Author information
Authors and Affiliations
Department of Biostatistics, Harvard T.H. Chan School of Public Health, Boston, Massachusetts
Carmen B. Rodriguez, Stephanie M. Wu, and Briana J.K. Stephenson
Division of Gynecologic Oncology, Department of Obstetrics and Gynecology and Reproductive Biology, Brigham and Women’s Hospital, Boston, Massachusetts
Harvard Medical School, Boston, Massachusetts
Stephanie Alimena
Dana-Farber Cancer Institute, Boston, Massachusetts
Carmen B. Rodriguez and Stephanie Alimena
Department of Health Policy and Management, Harvard T.H. Chan School of Public Health, Boston, Massachusetts
Alecia J. McGregor
Contributions
Corresponding author
Correspondence to Briana J.K. Stephenson.
Acknowledgments
@@
n(%) Profile 1 Profile 2 Profile 3 Profile 4 Profile 5 Profile 6 Profile 7 Profile 8
n=817 n= 426 n=236 n=285 n=196 n= 130 n=179 n=143
@@
Non-Hispanic White 776 (95.0) 275 (64.6) 213 (90.3) 271 (95.1) 170 (86.7) 94 (72.3) 169 (94.4) 135 (94.4)
Non-Hispanic Black 10 ( 1.2) 52 (12.2) 4 ( 1.7) 0 ( 0.0) 13 ( 6.6) 13 (10.0) 4 ( 2.2) 3 ( 2.1)
Hispanic 22 ( 2.7) 16 ( 3.8) 7 ( 3.0) 7 ( 2.5) 1 ( 0.5) 14 (10.8) 3 ( 1.7) 2 ( 1.4)
Other 7 ( 0.9) 78 (18.3) 11 ( 4.7) 6 ( 2.1) 9 ( 4.6) 6 ( 4.6) 1 ( 0.6) 1 ( 0.7)
@@
US-born 391 (47.9) 136 (31.9) 123 (52.1) 112 (39.3) 60 (30.6) 42 (32.3) 66 (36.9) 38 (26.6)
Foreign-born 373 (45.7) 195 (45.8) 98 (41.5) 158 (55.4) 112 (57.1) 59 (45.4) 106 (59.2) 100 (69.9)
Unknown 53 ( 6.5) 95 (22.3) 15 ( 6.4) 15 ( 5.3) 24 (12.2) 29 (22.3) 7 ( 3.9) 5 ( 3.5)
@@
Younger than 50 58 ( 7.1) 52 (12.2) 29 (12.3) 21 ( 7.4) 19 ( 9.7) 16 (12.3) 10 ( 5.6) 9 ( 6.3)
50-64 366 (44.8) 198 (46.5) 111 (47.0) 132 (46.3) 90 (45.9) 58 (44.6) 84 (46.9) 72 (50.3)
65 or older 393 (48.1) 176 (41.3) 96 (40.7) 132 (46.3) 87 (44.4) 56 (43.1) 85 (47.5) 62 (43.4)
@@
Private 405 (49.6) 130 (30.5) 118 (50.0) 136 (47.7) 84 (42.9) 62 (47.7) 70 (39.1) 59 (41.3)
Medicare 341 (41.7) 175 (41.1) 83 (35.2) 123 (43.2) 76 (38.8) 49 (37.7) 81 (45.3) 58 (40.6)
Public/Government 35 ( 4.3) 78 (18.3) 19 ( 8.1) 14 ( 4.9) 16 ( 8.2) 8 ( 6.2) 16 ( 8.9) 13 ( 9.1)
Other 22 ( 2.7) 39 ( 9.2) 13 ( 5.5) 9 ( 3.2) 18 ( 9.2) 8 ( 6.2) 10 ( 5.6) 12 ( 8.4)
Not insured 14 ( 1.7) 4 ( 0.9) 3 ( 1.3) 3 ( 1.1) 2 ( 1.0) 3 ( 2.3) 2 ( 1.1) 1 ( 0.7)
@@
None 24 ( 2.9) 18 ( 4.2) 9 ( 3.8) 13 ( 4.6) 8 ( 4.1) 6 ( 4.6) 7 ( 3.9) 6 ( 4.2)
Resection 792 (96.9) 403 (94.6) 227 (96.2) 269 (94.4) 188 (95.9) 124 (95.4) 169 (94.4) 137 (95.8)
Other Surgery/Unknown 1 ( 0.1) 5 ( 1.2) 0 ( 0.0) 3 ( 1.1) 0 ( 0.0) 0 ( 0.0) 3 ( 1.7) 0 ( 0.0)
@@
No radiation treatment 591 (72.3) 310 (72.8) 180 (76.3) 216 (75.8) 142 (72.4) 104 (80.0) 133 (74.3) 113 (79.0)
External beam radiation therapy (EBRT) 5 ( 0.6) 5 ( 1.2) 1 ( 0.4) 2 ( 0.7) 0 ( 0.0) 4 ( 3.1) 0 ( 0.0) 1 ( 0.7)
Brachytherapy 149 (18.2) 66 (15.5) 33 (14.0) 43 (15.1) 31 (15.8) 12 ( 9.2) 32 (17.9) 17 (11.9)
Other 72 ( 8.8) 45 (10.6) 22 ( 9.3) 24 ( 8.4) 23 (11.7) 10 ( 7.7) 14 ( 7.8) 12 ( 8.4)
@@
Stage I 687 (84.1) 345 (81.0) 186 (78.8) 246 (86.3) 164 (83.7) 105 (80.8) 151 (84.4) 119 (83.2)
Stage II 78 ( 9.5) 44 (10.3) 30 (12.7) 27 ( 9.5) 15 ( 7.7) 13 (10.0) 15 ( 8.4) 14 ( 9.8)
Stage III 27 ( 3.3) 12 ( 2.8) 11 ( 4.7) 5 ( 1.8) 10 ( 5.1) 5 ( 3.8) 2 ( 1.1) 4 ( 2.8)
Stage IV 25 ( 3.1) 25 ( 5.9) 9 ( 3.8) 7 ( 2.5) 7 ( 3.6) 7 ( 5.4) 11 ( 6.1) 6 ( 4.2)

### v2

@@ 3.2 Spatial Mapping of NSDoH
The MBMM profile assignments for each census tract were recorded and linked with MA census tract geographic data. We developed a web-based interactive map using Shiny to display the MBMM results, available at https://mhn38j-carmen-rodriguez.shinyapps.io/nsdoh_profiles_app_urban/. In this map, we also added an additional layer to provide further context to our MBMM defined profiles. Namely, the 2020 Urban Boundaries and Environmental Justice (EJ) Populations, obtained from the Massachusetts Bureau of Geographic Information Data Hub [https://gis.data.mass.gov]. The EJ data contains census block groups– subdivisions within census tracts–across the state that meet one or more of the following criteria: (i) the annual median household income is not more than 65 % of the statewide annual median household income; (ii) minorities comprise 40% or more of the population; (iii) 25 percent or more of households lack English language proficiency.
@@
Neighborhoods assigned to Profile 1 (advantaged NHW ) were found predominantly in surrounding areas outside of Boston (e.g., Berkshire, Hampshire, Essex, and Middlesex counties). Examining the overlaid urban boundaries reveals that most census tracts in this profile are located within urbanized areas. Additionally, the average proportion of non-Hispanic White individuals in these census tracts was 89%, consistent with our profile results describing social and community context indicators. Neighborhoods assigned to Profile 2 (disadvantaged racially/ethnically diverse (BHL+), more renter-occupied housing with limited EN proficiency) were identified in areas with large immigrant populations (e.g., Lawrence, Methuen, Lynn, Chelsea, Dorchester, Revere, and Springfield) and a high average proportion of Hispanic/Latino residents. Additionally, most block groups that met the criteria for EJ populations, specifically regarding limited English proficiency in households and minority status, were located within the census tracts included in this profile. Neighborhoods in Profile 3(working class lower educational attainment) were primarily in areas west of Boston (e.g., Franklin and Worcester County). Neighborhoods in Profiles 4 (racially/ethnically diverse (A+) and greater economic security and educational attainment) and 5 (racially/ethnically diverse (ABHL+), more renter-occupied housing with limited EN proficiency) were primarily in areas with a known university presence and immediately surrounding Boston (e.g., Cambridge, Allston, Brookline, Somerville). See Supplementary Table 1.2 in the Supplementary Materials Section 1 for more information about the census tracts’ population age and race-ethnicity distribution by NSDoH profile.
3.3 NSDoH Profiles and Receipt of Optimal Care for EC
Supplementary Table 2.1 in the Supplemental Materials Section 2 shows sociodemographic and clinical information of the analytical sample of EC patients with endometroid histology who received optimal care or NCCN guidelines adherent treatment (82.3%). These patients were predominantly non-Hispanic White (87.2%), foreign-born (49.7%), aged 50-64 (46.1%), privately insured (46.8%), diagnosed at stage I (91.5%), and with grade 1 tumors (52.9%). Patients were initially treated at large medical facilities (75.7%), academic medical centers (38.7%), and diagnosed by physicians specializing in family/internal medicine (47%).
Table 2 details the distribution of patient demographics based on the NSDoH profile at which they resided at the time of EC diagnosis. Race-ethnicity, birthplace, insurance status at diagnosis, and initial point of care facility characteristics were significantly associated with NSDoH profiles. For example, among patients residing in neighborhoods belonging to NSDoH Profile 2, 49% were foreign-born, 42.6% had Medicare or public insurance, and 37.6% were initially treated at academic medical centers.
Figure 2 summarizes the results of the regression analysis. Although not statistically significant, compared to patients who resided in neighborhoods belonging to the NSDoH Profile 1 (advantaged NHW), patients in the NSDoH Profile 2 (disadvantaged racially/ethnically diverse (BHL+), more renter-occupied housing with limited EN proficiency) had lower odds [OR = 0.80, 95% Credible Interval (0.58,1.11)] of receiving optimal care after adjusting for year of diagnosis, age at diagnosis, insurance status at diagnosis, and initial type of care facility.
We also examined the relationship between NSDoH profiles and the type of initial care facility (academic medical center versus other facilities). The results, adjusted for year, age, and insurance status at diagnosis, are detailed in Supplemental Materials Section 2. Compared to NSDoH Profile 1 (advantaged NHW), patients who resided in NSDoH Profiles 4 (racially/ethnically diverse (A+) and greater economic security and educational attainment) and 5 (racially/ethnically diverse (ABHL+), more renter-occupied housing with limited EN proficiency ) were significantly associated with higher odds of receiving care at an academic medical center, while those in NSDoH Profile 3 (working class lower educational attainment) had lower odds, though not statistically significant.
Figure 2: Adjusted odds of receiving optimal care by NSDoH profiles (Ref= NSDoH Profile 1; advantaged NHW) for patients diagnosed with EC from 2015-2019. Bars show 95% Credible Intervals. Shorthand names for NSDoH Profiles: 1) advantaged NHW, 2) disadvantaged racially/ethnically diverse (BHL+), more renter-occupied housing with limited EN proficiency, 3) working class lower educational attainment, 4) racially/ethnically diverse (A+) and greater economic security and educational attainment, 5) racially/ethnically diverse (ABHL+), more renter-occupied housing with limited EN proficiency.
4 Discussion
We used data from the ACS to derive NSDoH profiles using a fully Bayesian multivariate Bernoulli mixture model. To our knowledge, this is the first study to apply a Bayesian mixture model to characterize neighborhood-level data in this context. This approach allowed us to estimate the appropriate number of profiles directly from the data, eliminating the need for post-hoc testing. Our model identified five NSDoH profiles, which we characterized using neighborhood SDoH variables belonging to four thematic domains: household conditions and resources, economic security, educational attainment, and social and community context.
We used these profiles to conduct a regression analysis to examine the association of NSDoH profiles and optimal care for EC. Our findings, although not statistically significant, suggest some neighborhood profiles experience lower odds of optimal EC care, compared to patients in the advantaged NH White profile (Profile 1). Specifically, patients residing in neighborhood profiles with higher proportions of racially/ethnically diverse residents and greater housing and educational burden had lower odds of receiving optimal care. These findings are comparable to Rodriguez et al. where they examined the association between neighborhood socioeconomic status (NSES) using the Yost Index and adherence to the NCCN guidelines from 2006-2015 [12]. Consistent with our findings, patients residing in the most disadvantaged neighborhoods with the lowest NSES had lower odds of receiving NCCN guideline-concordant care compared to those in the highest NSES group. In their study, 59.5% of patients received treatment adhering to NCCN guidelines, with the lowest adherence seen among Black, Latina, and American Indian/Alaska Native women (57.1%, 54.5%, and 52.7%, respectively). These proportions are lower than those in our study (82.3%) but comparable to a study using the Women’s Health Initiative (WHI), where they found 80% of patients received NCCN adherent treatment for EC [49]. Our results focused specifically on Massachusetts, while their analysis covered the entire United States. Additionally, similar to the Women’s Health Initiative, our patient population was predominantly non-Hispanic White and limited to individuals engaged in the healthcare system. This demographic composition may further obscure disparities in access to and receipt of guideline-concordant care. Systemic barriers such as insurance status, provider bias, and structural racism have contributed to lower levels of healthcare engagement among some racial and ethnic minority populations, resulting in delayed screening, diagnosis, and more advanced disease at presentation [50]. This is also important, as differences in standard treatment regimens may exist for some groups due to variations in the histologic subtypes and stages of endometrial cancer they present with [51]. Lastly, beyond the limitations of retrospective observational studies, the Massachusetts Cancer Registry lacks information on comorbidities, which may also impact differences in treatment adherence and recommendations. This could also explain the lack of guideline-concordant treatment rates for certain demographics.
Several studies have relied on the Yost index as a reliable measure of socioeconomic status at the neighborhood level [12, 9, 13, 52]. This measure of neighborhood disadvantage uses a subset of our variables, which yields a different narrative of SES that is not necessarily reflective of the socioeconomic barriers of healthcare access. The Yost index is derived via a one component PCA. This approach is a powerful and widely used method to reduce the dimensionality of high-dimensional datasets, but it suffers from interpretability issues because it outputs a linear combination of the original indicators [53]. These results, by focusing on a composite score, could mask multidimensional deprivation patterns across neighborhoods. Our approach, which relied on a mixture model framework, provides a holistic way to identify patterns of multiple interrelated exposures jointly. This approach provides interpretable clusters of neighborhoods, making it easier to understand the heterogeneity of neighborhood data, as well as the defining characteristics of each profile. The Bayesian framework of the MBMM provides the flexibility of estimating the model parameters using prior information, including the parameter describing the number of profiles [44, 45, 54]. Utilizing a Bayesian framework allowed us to borrow information within and across other neighborhoods to improve the precision of our estimates, which improved the identifiability of our profiles.
This model is still met with limitations. First, model-based clustering is reflective of the data we use. In this study, we used 5-year estimates from the ACS from 2015-2019 for MA. Analysis using different ACS survey waves, different SDoH variables, or different geographies may yield different NSDoH profiles. For example, our results were based on binarized thresholds defined by median values for the state of Massachusetts. A different state would yield different cutoffs and, ultimately, different profiles. Second, the decision to dichotomize the SDoH variables was based on the skewed distributions found in the 2015-2019 cycles. This can sometimes result in a slight loss of information and reduced precision of the profile estimates since the full distribution of these variables was not considered. Future work should explore the development and implementation of a fully Bayesian multivariate beta mixture model that can flexibly accommodate bounded data, such as what we see in census-level data. Third, our analysis used aggregate data from the Census to construct NSDoH profiles. These variables are reflective of population-level characteristics rather than individual-level data. To avoid ecological fallacies, relationships observed at the aggregate level cannot be assumed to hold for individual EC patients residing within these neighborhoods. Lastly, consistent with other similar studies, our analysis is cross-sectional [23] and does not account for the changes in neighborhood demographics over time as a result of fair housing policies and gentrification. We focused our time frame on 2015-2019, which was based on the ACS 5-year estimates of data. Future studies should incorporate multiple survey cycles to assess how demographic shifts influence NSDoH profiles and, consequently, the access to and quality of healthcare and other outcomes for their residents.
Bayesian mixture models offer promising applications for neighborhood-level data. Our approach enabled us to characterize and assign NSDoH profile patterns to neighborhoods in Massachusetts. Geospatial mapping of NSDoH profiles demonstrated how we can leverage these tools to identify areas for targeted interventions. Using our NSDoH profiles to assess association with health outcomes, such as receipt of optimal care for EC, may give a more nuanced understanding of how SDoH overlap and co-occur within communities rather than in isolation to shape health experiences and outcomes. While this study focused on receipt of optimal care for EC patients, the derived profiles are translatable to a myriad of other outcomes and exploratory analyses.
Competing interests
The authors declare that they have no competing interests.
Acknowledgments
Funding
This project was supported by the Harvard Data Science Initiative Bias2 award given to the Stephenson Lab. Dr. Stephenson is supported by NHLBI 1K01HL166442. Carmen B. Rodríguez is supported by NIH T32CA009337. The funders had no role in the conceptualization, design, data analysis, decision to publish, or preparation of the manuscript.
Author contributions statement
Data availability statement
The R code used for this manuscript has been made publicly available in GitHub, https://github.com/cbrodriguez01/ecbayesbinmix. This also includes all the American Community Survey data we used for the multivariate Bernoulli mixture model, which is also publicly available at https://www.census.gov/programs-surveys/acs/data.html. The data that support the findings from the regression analysis in this study are available from the Massachusetts Cancer Registry-Massachusetts Department of Public Health (MCR-MDPH), but restrictions apply to the availability of MCR data, which were used under license for the current study, and so are not publicly available. Data are, however, available from the authors upon reasonable request and with permission of MCR-MDPH.
@@ Acknowledgments
Table 1: Median estimate for Massachusetts for selected neighborhood social determinants of health and ACS 2015-2019 table identification.
NSDoH ACS Variable (ACS Table) Median
@@
% Crowding in household (DP04$\dagger$ 1.28
% Occupied housing units without complete plumbing (DP04$\ddagger$ 0
Estimate median household income in the past 12 months (inflation-adjusted; B19013)§$\S$ 82265
@@
% Population aged 25 years or older with no high school diploma (DP02) 6.5
% Language other than English: Speak English less than “very well”(DP02) 5.8
@@
  • $\dagger$
    Occupied housing units with 1.01 to 1.50 and 1.51 or more occupants per room/ All occupied housing units for the same calendar year[55].
  • $\ddagger$
    We binarized this variable differently because the median was 0. Therefore, census tracts for which there were no households lacking complete plumbing were categorized as “0”, and the remaining tracts were categorized as “1” to indicate lack of plumbing.
  • §$\S$
    Census tracts for which the median household income was $\geq$ to the state median were categorized as ”0”, and remaining tracts were categorized as ”1” to indicate lower income.
@@
n(%) Profile 1 Profile 2 Profile 3 Profile 4 Profile 5
n= 952 n= 439 n=494 n=340 n=187
@@
Non-Hispanic White 910 (95.6) 286 (65.1) 459 (92.9) 313 (92.1) 135 (72.2)
Non-Hispanic Black 7 ( 0.7) 54 (12.3) 12 ( 2.4) 6 ( 1.8) 20 (10.7)
Hispanic 21 ( 2.2) 13 ( 3.0) 8 ( 1.6) 13 ( 3.8) 17 ( 9.1)
Other 11 ( 1.2) 82 (18.7) 9 ( 1.8) 7 ( 2.1) 10 ( 5.3)
@@
US-born 439 (46.1) 124 (28.2) 166 (33.6) 171 (50.3) 68 (36.4)
Foreign-born 458 (48.1) 217 (49.4) 298 (60.3) 144 (42.4) 84 (44.9)
Unknown 55 (5.8) 98 (22.3) 30 ( 6.1) 25 ( 7.4) 35 (18.7)
@@
Younger than 50 69 ( 7.2) 51 (11.6) 39 ( 7.9) 32 ( 9.4) 23 (12.3)
50-64 433 (45.5) 202 (46.0) 236 (47.8) 151 (44.4) 89 (47.6)
65 or older 450 (47.3) 186 (42.4) 219 (44.3) 157 (46.2) 75 (40.1)
@@
Private 463 (48.6) 132 (30.1) 204 (41.3) 177 (52.1) 88 (47.1)
Medicare 403 (42.3) 187 (42.6) 203 (41.1) 126 (37.1) 67 (35.8)
Public/Government 46 ( 4.8) 78 (17.8) 44 ( 8.9) 19 ( 5.6) 12 ( 6.4)
Other 27 ( 2.8) 38 ( 8.7) 38 ( 7.7) 12 ( 3.5) 16 ( 8.6)
Not insured 13 ( 1.4) 4 ( 0.9) 5 ( 1.0) 6 ( 1.8) 4 ( 2.1)
@@
None 28 ( 2.9) 19 ( 4.3) 20 ( 4.0) 13 ( 3.8) 11 ( 5.9)
Resection 920 (96.6) 415 (94.5) 471 (95.3) 327 (96.2) 176 (94.1)
Other Surgery/Unknown 4 ( 0.4) 5 ( 1.1) 3 ( 0.6) 0 ( 0.0) 0 ( 0.0)
@@
No radiation treatment 691 (72.6) 328 (74.7) 367 (74.3) 260 (76.5) 143 (76.5)
External beam radiation therapy (EBRT) 4 ( 0.4) 4 ( 0.9) 2 ( 0.4) 3 ( 0.9) 5 ( 2.7)
Brachytherapy 174 (18.3) 65 (14.8) 80 (16.2) 46 (13.5) 18 ( 9.6)
Other 83 ( 8.7) 42 ( 9.6) 45 ( 9.1) 31 ( 9.1) 21 (11.2)
@@
Stage I 806 (84.7) 361 (82.2) 416 (84.2) 278 (81.8) 142 (75.9)
Stage II 88 ( 9.2) 42 ( 9.6) 42 ( 8.5) 39 (11.5) 25 (13.4)
Stage III 32 ( 3.4) 11 ( 2.5) 13 ( 2.6) 11 ( 3.2) 9 ( 4.8)
Stage IV 26 ( 2.7) 25 ( 5.7) 23 ( 4.7) 12 ( 3.5) 11 ( 5.9)